If cos x = 0.42, which of the following is cos(x +12π)?

A. 0.042
B. 0.04
C. 0.54
D. 0.42
1
Expert's answer
2013-06-06T10:24:56-0400
Cosine function is periodic with period 2pi.
Due to it, cos(x+2pi*k)=cos x for any integer k.
In our case k=6: cos(x+12pi)=cos(x+6*2pi)=cos x
Thus, correct answer is D.